OpenStudy (anonymous):

2^(5/2)-2^(3/2) Why is the answer to this equation 2^(3/2)

OpenStudy (asnaseer):

think of how \(2^{5/2}\) can be rearranged

OpenStudy (asnaseer):

\[2^{5/2}=2^{(3+2)/2}=2^{3/2+2/2}=2^{3/2+1}=2\times2^{3/2}\]

OpenStudy (anonymous):

@GluttonousIntention09 this is why: \[2^{2.5} - 2^{1.5}\] \[= 2^{(2+0.5)} - 2^{(1+0.5)}\] =\[= (2^{2} * 2^{0.5}) - (2^{1} * 2^{0.5)}\] factor out the 2^0.5 \[= 2^{0.5}(2^{2} - 2^{1)}\] \[=2^{0.5} * 2^{1}\] \[=2^{1.5} = 2^{3/2}\]